OpenStudy (anonymous):

(5x2^1/2+2)(2x6^1/2+1)

OpenStudy (wolfboy):

ORDER OF OPERATIONS!!!!!!! PEMDAS

OpenStudy (wolfboy):

For the first pair of parenthisis 5*2=10/2=7

OpenStudy (wolfboy):

For the second pair 2*6=12/2=6+1=7

OpenStudy (wolfboy):

7*7=49